Part 1: Prove that the system of linear congruences in one variable
x ≡ b1 mod m1
x ≡ b2 mod m2
is solvable if and only if (m1, m2)|b1 − b2. In this case prove that the solution is unique modulo [m1, m2].
Hint: follow the proof of the Chinese Remainder Theorem.
Part 2: Solve the system of congruences
x ≡ 3 mod 4
x ≡ 1 mod 6
1
Expert's answer
2019-10-18T13:27:34-0400

If the given system has a solution, then

"x ≡ b_1\ mod\ m_1,\ x ≡ b_2\ mod\ m_2,"

therefore

"x=b_1+k_1m_1,\ x=b_2+k_2m_2" for some integers "k_1\ and\ k_2"

"b_1-b_2=k_2m_2-k_1m_1"

"(m_1,m_2)" divides "m_1" and "m_2" "\implies"

"b_1-b_2=k_2l_2(m_1,m_2)-k_1l_1(m_1,m_2)" for some integers "l_1\ and \ l_2"

"b_1-b_2=(m_1,m_2)(k_2l_2-k_1l_1)"

this means "(m_1,m_2)|b_1-b_2" .

\

If "(m_1,m_2)|b_1-b_2" ,then

"b_1-b_2=k(m_1,m_2)"

Using Bezout's identity

"(m_1,m_2)=am_1+bm_2" for some integers "a\ and\ b"

"b_1-b_2=k(am_1+bm_2)\implies"

"b_1-kam_1=b_2+kbm_2"

Now, if we set "x=b_1-kam_1=b_2+kbm_2", then "x" is the solution of the given congruence.

If "x_1\ and \ x_2" are two solutions of the given congruence, then

"x_1=b_1+k_1m_1,\ x_2=b_2+k_2m_2\implies"

"x_1-x_2=b_1-b_2+k_1m_1-k_2m_2"

"(m_1,m_2)" divides "m_1\ and\ m_2" and we proved that "(m_1,m_2)|(b_1-b_2)", therefore

"x_1-x_2" is divisible by "(m_1,m_2)"

and the solution is unique modulo "(m_1,m_2)."

\

x ≡ 3 mod 4

x ≡ 1 mod 6

3x ≡ 9 mod 12

2x ≡ 2 mod 12

3x-2x=x≡7 mod 12

x=7 is the solution of the given congruence.

Answer: x=7.